OpenStudy (ineedhelpfast12):

solve the linear system using substitution. x=y+3 2x-y=5

pooja195 (pooja195):

x=y+3 <--equation 1 2x-y=5 <---equation 2 x value is given so plug it into equation 2 \[\huge~\rm~\bf~2(y+3)−y=5 \]

pooja195 (pooja195):

Solve for y.

TheSmartOne (thesmartone):

Plug in the first equation into the second equation. Can you tell me what you get? :)

TheSmartOne (thesmartone):

dang the lag :/

OpenStudy (ineedhelpfast12):

 y=-1